4. Problems and Applications Q4 An economy consists of three workers: Rajiv, Yakov, and Charles. Each works 10 hours a day and can produce two services: mowing lawns and washing cars. In an hour, Rajiv can either mow 2 lawns or wash 1 car; Yakov can either mow 1 lawn or wash 1 car; and Charles can either mow 1 lawn or wash 2 cars. For each of the scenarios listed in the following table, determine how many lawns will be mowed and how many cars will be washed per day and enter these values into the corresponding row Scenario Lawns Mowed Cars Washed All three spend all their time mowing lawns. (A) All three spend all their time washing cars(B) All three spend half their time on each activity. (C) Rajiv spends half his time on each activity, while Yakov only washes cars and Charles only mows lawns. (D) In the following table, identify the opportunity cost of washing cars for each worker. Worker Opportunity Cost of Washing Cars Yakov CharlesAssume that the resources best suited to producing a particular service are preferentially used in the production of that service and that as the economy moves down along the production possibilities frontier, one worker at a time is transferred from mowing lawns to washing cars. Using the blue points (circle symbol), graph the production possibilities frontier (PPF) for this economy on the following graph. Then use the black point (plus symbol) to identify point A, the green point (triangle symbol) to identify point B, the orange point (square symbol) to identify point C, and the purple point (diamond symbol) to identify point D on the graph 50 45 PPF 40 E 30 25 20 15 10 0 5 10 15 20 25 30 35 4 45 50 Quantity of Cars Washed True or False: The production possibilities frontier consists of straight-line segments, rather than being smoothly bowed outward throughout, because each worker faces a constant trade-off between mowing lawns and washing cars. O True FalseTrue or False: The production possibilities frontier consists of straight-line segments, rather than being smoothly bowed outward throughout, because each worker faces a constant trade-off between mowing lawns and washing cars O True O False Indicate whether each of the following allocations is efficient or inefficient. Allocation Efficient Inefficient6. Problems and Applications Q6 Classify each of the following statements as positive or normative Positive Normative Statement Society faces a short-run trade-off between inflation and unemployment. The government should provide basic healthcare to all citizens Limiting unemployment benefits to only two years increases the incentive to look for jobs.O The government should limit unemployment benefits to last for only two years. A price set at lower than the equilibrium price will lead to shortages. Grade It Now Save & Continue Continue without saving
Answer #1

4. (A) Lawn mowed by Rajiv, Yakov and Charles = 2 x 10hours + 1 x 10 + 1 x 10 = 20 + 10 + 10 = 40

(B) Car washed by all = 1 x 10 + 1 x 10 + 2 x 10 = 40

(C) Lawn mowed = 2 x 5 hours + 1 x 5 + 1 x 5 = 10 + 5 + 5 = 20

Car washed = 1 x 5 hours + 1 x 5 + 2 x 5 = 20

(D) Lawn mowed = 2 x 5 hours + 1 x 10 hours = 20

Car washed = 1 x 5 + 1 x 10 = 15

Opportunity cost is the value of next best alternative foregone.

1. Rajiv:

1 car wash = 2 lawn mowed

It implies opportunity cost of washing car is 2 Lawn mowed.

2. Yakov:

1 car wash = 1 lawn mowed

Opportunity cost of 1 car wash is 1 lawn mowed.

3. Charles:

2 car wash = 1 lawn mowed

1 car wash = 1/2 or 0.5 lawn mowed

Opportunity cost is 0.5 lawn mowed.

50 T PPF 40 E 30 20 0 15 10 T 0 5 15 20 25 03540 45 50 Quantity of Cars Washed

True.

A, B, C - These are efficient because lie on the PPF.

D is inefficient because it lies below PPF which kept some resources idle.